OpenStudy (anonymous):

PLEASE HELP. multiply or divide. show your work. 1. 5x+10/x+2*2x/4x-10 2. 3n^2-n/n^2-1 ÷ n^2/n+1 3. the width W of a rectangular swimming pool is 7x^2. the area A of the pool is 7x^3-42x^2. what is an expression for the length of the pool?

OpenStudy (anonymous):

5x+10/x+2*2x/4x-10 6x+10 8x+10 12x+10/2/-10 12x+5/-10 12x+-2 -12 -12 x = 24 I think

OpenStudy (anonymous):

whatta bout the other ones?

OpenStudy (jhannybean):

1.\[\frac{5x+10}{x+2} \cdot \frac{2x}{4x-10}\]\[\frac{5\color{red}{(x+2)}}{\color{red}{x+2}}\cdot \frac{2x}{4x-10}\]\[=\frac{5\cdot 2x}{4x-10}\] Simplify.

OpenStudy (anonymous):

would that be the asnwer? im confused..

OpenStudy (jhannybean):

2.\[\frac{3n^2-n}{n^2-1} \div \frac{n^2}{n+1} \implies \frac{3n^2-n}{n^2-1} \cdot \frac{n+1}{n^2} \]Factor out like terms from each fraction. \[\frac{3\color{red}{n^2}-\color{red}{n}}{\color{blue}{n^2-1}} \cdot \frac{\color{blue}{n+1}}{\color{red}{n^2}}\]\[\color{blue}{a^2-b^2 = (a-b)(a+b)}\]Simplify.

OpenStudy (jhannybean):

I'm not giving you answers, I'm just helping you work them out.

OpenStudy (jhannybean):

You have to simplify them yourself and find the answer.

OpenStudy (anonymous):

idk how to do that

OpenStudy (anonymous):

honestly, i've never been good at math. i cant comprehend it good

OpenStudy (anonymous):

doesn't she need to divide the or subtract the fore from 4x and 10

OpenStudy (jhannybean):

\[4x-10 \rightarrow 4\left(x-\frac{10}{4}\right)\]

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

3) question area =wl=7x^2(l) =7x^3-42x^2 given l ength =7x^3-42x^2 --------------- 7x^2 that's a division simply divide by 7 and x^2 gives the answer